If 60 gallons of a 20% solution is needed and all that's available is a 10% solution and a40% solution how many of each is needed
A

Answer by mananth(16946)

percent ---------------- quantity
Solution I 10.00% ---------------- x gallons
solution II 40.00% ------ 60 - x gallons
Mixture 20.00% ---------------- 60
Total 60 gallons
10.00% x + 40.00% ( 60 - x ) = 20.00% * 60
10 x + 40 ( 60 - x ) = 1200
10 x + 2400 - 40 x = 1200
10 x - 40 x = 1200 - -2400
-30 x = -1200
/ -30
x = 40 gallons 10.00% Solution I
20 gallons 40.00% solution II

m.ananth@hotmail.ca